OpenStudy (anonymous):

80 more than twice a number is 134. What is the number?

OpenStudy (hoblos):

2n+80 = 134

OpenStudy (anonymous):

2x+80 =134

OpenStudy (anonymous):

2c+80=134 ,c=(134_80)\2

OpenStudy (anonymous):

27

OpenStudy (hoblos):

btw n , x and c are the same :P